OpenStudy (anonymous):

What is the graph of the system y = -2x + 3 and 2x + 4y = 8?

pooja195 (pooja195):

we need to find the coordinates what method would you like to use?

pooja195 (pooja195):

Elimination Substitution Graphing

pooja195 (pooja195):

@TheMathematiciansAnvi

OpenStudy (anonymous):

I prefer elimination

pooja195 (pooja195):

\[\frac{ y = -2x + 3 }{ 2x + 4y = 8? }\] you want to eliminate either the X or Y what number would you multiply by?

pooja195 (pooja195):

ignore the ? mark next to the 8

pooja195 (pooja195):

wait

pooja195 (pooja195):

Substitution would be easier here

OpenStudy (anonymous):

I'm sorry, I have a question, do I keep the bottom equation in standard form or do I change it to slope before eliminating?

pooja195 (pooja195):

none we should do substitution since it already tells u what y equals

OpenStudy (anonymous):

Okay ya that would be smart

OpenStudy (anonymous):

It would be 2x-8x+12=8

OpenStudy (anonymous):

so -6x+12=8

OpenStudy (anonymous):

Is that so

pooja195 (pooja195):

yes now finish solving it :)

OpenStudy (anonymous):

x= 4/6?

pooja195 (pooja195):

hmm i think thats right https://www.desmos.com/calculator <----- that should help too

OpenStudy (anonymous):

Okay, so now that I have x=4/6 I just substitute x in the equation to find y?

pooja195 (pooja195):

yep

OpenStudy (anonymous):

Okay thanks, I have one more. I don't really understand the graphing with linear equations.

pooja195 (pooja195):

ok :) i can help report it also use this https://www.desmos.com/calculator it will graph it for you and you will only need to find the coordinates

OpenStudy (anonymous):

3x + 4 = 5x - 10 3x - 4 = 5x - 10 3x + 4 = 5x + 10 3x - 4 = 5x + 10

OpenStudy (anonymous):

BTW, I already graphed the last one I just wanted to get a bit ahead

pooja195 (pooja195):

LOL ok good :) ok so here im not sure but i know you can definetly find out by plugging the coordinate to each of the options

OpenStudy (anonymous):

Okay, I always try to avoid that in order to know how to do it but I guess I could use a little shortcut :P

OpenStudy (anonymous):

So basically, once I plug-in the x-coordinate into each equation whichever I get that both sides are equal is my answer, am I correct?

OpenStudy (anonymous):

as in 1=1 would be correct and 2=1 wouldn't

pooja195 (pooja195):

yes so if 0=0 <-------- many solutions 0=9<---------no solution

OpenStudy (anonymous):

Okay good

OpenStudy (anonymous):

@pooja195 Thanks for the help!